Use subsitutuion to determine wether the point is on the line.
y= x-4; A (5,1)
1
Expert's answer
2012-09-11T09:09:56-0400
y= x-4
Substitute x=5, y=1:
1=5-4
1=1 that's right therefore given point A is on the line y=x-4